What is the molecular weight of a nonvolatile, covalent compound of 7.58 g of it in 100.0 g of water yields a solution that boils at 100.21 degree Celsius at laboratory pressure? Pure water has a boiling point of 99.88 degree Celsius at the same pressure
1
Expert's answer
2020-12-17T07:43:29-0500

∆T=Kbm.

m=∆T/Kb. = 100.21-99.88/0.512 = 0.64453125.

n= 0.64453125 x 0.1 = 0.064453125 moles.

Mr= m/n = 7.58/0.64453125 = 117,604848.

If this molecule two ions 117,604848/2 = 58.8

And it is NaCl.